Average Rate of Return

The following data are accumulated by Lone Peak Inc. in evaluating two competing capital investment proposals:

3D Printer Truck
Amount of investment $60,000 $28,000
Useful life 4 years 9 years
Estimated residual value 0 0
Estimated total income over the useful life $7,800 $11,340

Determine the expected average rate of return for each proposal. If required, round your answers to one decimal place.

3D Printer %
Truck %
Answer #1
Average income:
3D Printer 1950 =7800/4
Truck 1260 =11340/9
Average investment:
3D Printer 30000 =(60000+0)/2
Truck 14000 =(28000+0)/2
Average rate of return:
3D Printer 6.5% =1950/30000
Truck 9.0% =1260/14000